# Searching for light long-lived particles at SHiNESS

### v2

@@ 4.1 Light binos
In Fig. 5, we show two sensitivity plots in the plane $\lambda_{312}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$ and $\lambda_{312}/m^{2}_{\tilde{f}}=\lambda_{131}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$, respectively. In the upper plot, we fix the light neutralino’s mass at 10 MeV, 20 MeV, 60 MeV, and 100 MeV. The present bounds on the RPV couplings for $m_{\tilde{f}}=1$ TeV and 2 TeV are shown, and we find that in the optimal case ($m_{\tilde{\chi}^{0}_{1}}=60$ MeV), new parameter space up to more than two orders of magnitude beyond the current bounds can be excluded by SHiNESS in 2 years. The lower plot of Fig. 5 indicates that SHiNESS is predicted to probe unexcluded values of $\lambda_{312}/m^{2}_{\tilde{f}}=\lambda_{131}/m^{2}_{\tilde{f}}$ more than one order of magnitude beyond the present limits in two years, for $m_{\tilde{\chi}^{0}_{1}}$ between 10 MeV and 100 MeV.
@@ 4.2 Weak-violating ALPs coupled to electrons
The projected sensitivities of SHiNESS with 2-year data collation are presented in Fig. 6, in the plane $c_{ee}/\Lambda$ vs. $m_{a}$. The current bounds stemming from electron beam-dump experiments Konaka:1986cb ; Riordan:1987aw ; Bjorken:1988as ; Bross:1989mp , supernova SN1987A Lucente:2021hbp , and searches for flavor-changing-neutral-current decays of mesons KTeV:2003sls ; LHCb:2015ycz ; NA62:2020xlg , are shown in gray. Further, bounds obtained in searches for leptonic decays of charged mesons are displayed in blue for $\pi^{+}$ decays SINDRUM:1986klz , in green for $K^{+}$ decays Poblaguev:2002ug , and in brown for the CHARM experiment CHARM:1985anb . Particularly, rough estimates for potential bounds obtained from charged-pion and kaon decays are also overlaid with dashed lines, for the PIONEER experiment PIONEER:2022yag (blue), kaon factories Goudzovski:2022vbt (green), and the LHC (purple) Altmannshofer:2022ckw . We observe that while the potential of these future experiments is estimated to probe mainly the unexcluded parameter region of $c_{ee}/\Lambda\gtrsim 6\times 10^{-4}$ GeV-1, SHiNESS can be sensitive to a unique, allowed region $c_{ee}/\Lambda\sim\mathcal{O}(10^{-5})$ GeV-1 for $m_{a}$ between roughly 10 MeV and 120 MeV.
Figure 6: The projected sensitivity of SHiNESS with 2 years’ operation to the weak-violating electrophilic ALP, displayed in the $(m_{a},c_{ee}/\Lambda)$ plane. The area inside the red curve can be excluded by SHiNESS. The gray area in the background, as well as the blue, green, and brown shaded regions, are the currently excluded parameter space, reproduced from Ref. Altmannshofer:2022ckw ; see the text for detail. Ref. Altmannshofer:2022ckw also provided estimate of potential sensitivities from the PIONEER experiment PIONEER:2022yag (dashed blue), kaon factories Goudzovski:2022vbt (dashed green), as well as a benchmark of the LHC (dashed purple) Altmannshofer:2022ckw . See also the discussion in Sec. 2.2.
@@ 5 Conclusions
In this work, we have investigated the potential of the SHiNESS experiment proposed to be installed at the ESS facility, to search for light long-lived particles. Following the SHiNESS proposal Soleti:2023hlr , we focus on the signature of LLPs at SHiNESS as excesses in electron-positron pair events, and selection cuts on the final-state leptons are imposed in order to suppress background events to an acceptable level. Moreover, in order to estimate the cut efficiencies and acceptance of SHiNESS to LLPs, we have performed numerical simulations.
The focus of this work is on LLPs produced in rare decays at rest of $\pi^{+}$ mesons that arise from the 2-GeV proton beam at the facility hitting a rotating tungsten target and lose energy as they propagate. Specifically, we choose two theoretical models predicting light, MeV-scale LLPs, namely, a light bino in the RPV-SUSY, and a weak-violating electrophilic ALP. We consider production modes of the light bino including decays of $\pi^{+}\to e^{+}\tilde{\chi}^{0}_{1}$ and $\mu^{+}\to e^{+}\overset{(-)}{\nu}\tilde{\chi}^{0}_{1}$, where $\mu^{+}$ stems from the $\pi^{+}\to\mu^{+}\nu_{\mu}$ decay at rest and also decays at rest immediately. The light bino then travels and decays inside SHiNESS into an electron-positron pair in association with a neutrino. Various benchmark scenarios, each with two non-vanishing RPV couplings mediating the production and decay of the light bino, respectively, are studied. Numerical results are then obtained and presented in the plane of either the production coupling vs. the decay coupling, or the RPV couplings (assumed to be equal) versus the light bino’s mass. While for some benchmark scenarios we find SHiNESS is unable to probe unexcluded parameter space of the model for reasons of kinematics or too strong current bounds, for other benchmarks SHiNESS is predicted to be able to exclude large, new territories in the parameter space with just two years’ data collection.
For the ALP, we focus on the $\pi^{+}\to e^{+}\nu_{e}a$ decay for its production and on $a\to e^{-}e^{+}$ for its signal decay. We assume a single, weak-violating term in an effective Lagrangian with the ALP coupled to a pair of electrons, that, after integration by parts, leads to both a helicity-suppressed three-point-vertex term, and a helicity-unsuppressed four-point-vertex term. The latter contribution, involving in the vertex an ALP, an electron, an electron neutrino, as well as a $W$-boson, results in large production rates of the ALP from the $\pi^{+}\to e^{+}\nu_{e}a$ decay. In the numerical results, we find that with solely two years’ operation, SHiNESS can probe a unique, unexcluded part of the parameter space of $c_{ee}/\Lambda\sim\mathcal{O}(10^{-5})$ GeV-1 for $m_{a}$ between 10 MeV and 120 MeV, complementary to other future experiments which would be sensitive to larger values of $c_{ee}/\Lambda$ in a similar range of $m_{a}$.
@@
We conclude that with the ESS user program planned to begin in 2025, to implement a detector such as SHiNESS at the unallocated space in the D03 experimental hall is timely and has an exciting prospect for searching for BSM physics such as LLPs. We believe our work with the positive results highly motivates the construction and operation of the SHiNESS experiment.
@@ Acknowledgements
We thank Liangwen Chen and Martin Schürmann for useful discussions. Y.Z. is supported by the National Natural Science Foundation of China under Grant No. 11805001 and the Fundamental Research Funds for the Central Universities under Grant No. JZ2023HGTB0222. W.L. is supported by National Natural Science Foundation of China (Grant No. 12205153).
Appendix A Computation procedure of $N_{S}^{90}$
We explain the computation procedure of the number of NP signal events $N_{S}$ required for estimating the sensitivity reach of the proposed SHiNESS experiment at 90% C.L., with an unbinned $\chi^{2}$ test in a null hypothesis.
As mentioned in Sec. 3, the background-event number for the electron-positron pair search is expected to be $61\pm 31$ per year after all event-selection requirements such as $E_{e^{-}},E_{e^{+}}>17$ MeV, are imposed Soleti:2023hlr . We thus estimate that for two years SHiNESS should suffer from $N_{B}=61\times 2=122$ background events with an uncertainty of $\sigma_{N_{B}}=\sqrt{2}\times 31\approx 43.84$ events. For the $\chi^{2}$ distribution with one degree of freedom as is the case of this unbinned test, a critical value $\chi^{2}_{\text{c.v.}}$ of 2.71 is required for the 90% C.L.
To compute the $\chi^{2}$ value we use the following expression,
$\chi^{2}=N_{S}^{2}/\sigma^{2}_{N_{B}},$ (23)
which is required to be larger than or equal to $\chi^{2}_{\text{c.v.}}$, in order to reject the null hypothesis, implying $N_{S}\geq\sqrt{\chi^{2}_{\text{c.v.}}}\cdot\sigma_{N_{B}}\approx 72.17$ corresponding to $N_{S}^{90}$.

### v3

@@ 4.1 Light binos
In Fig. 5, we show two sensitivity plots in the $\lambda_{312}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$ and $\lambda_{312}/m^{2}_{\tilde{f}}=\lambda_{131}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$ planes, respectively. In the left plot, we fix the light neutralino’s mass at 10 MeV, 20 MeV, 60 MeV, and 100 MeV. The present bounds on the RPV couplings for $m_{\tilde{f}}=1$ TeV and 2 TeV are shown, and we find that in the optimal case ($m_{\tilde{\chi}^{0}_{1}}=60$ MeV), new parameter space up to more than two orders of magnitude beyond the current bounds can be excluded by SHiNESS with two years’ data collection. The right plot of Fig. 5 indicates that SHiNESS is predicted to probe unexcluded values of $\lambda_{312}/m^{2}_{\tilde{f}}=\lambda_{131}/m^{2}_{\tilde{f}}$ more than one order of magnitude beyond the present limits in two years’ operation, for $m_{\tilde{\chi}^{0}_{1}}$ between 10 MeV and 100 MeV.
@@ 4.2 Weak-violating ALPs coupled to electrons
The projected sensitivities of SHiNESS with two-year data collation are presented in Fig. 6, in the plane $c_{ee}/\Lambda$ vs. $m_{a}$. The current bounds stemming from electron beam-dump experiments Konaka:1986cb ; Riordan:1987aw ; Bjorken:1988as ; Bross:1989mp , supernova SN1987A Lucente:2021hbp , and searches for flavor-changing-neutral-current decays of mesons KTeV:2003sls ; LHCb:2015ycz ; NA62:2020xlg , are shown in gray. Further, bounds obtained in searches for leptonic decays of charged mesons are included in the gray area, taking into account $\pi^{+}$ decays SINDRUM:1986klz , $K^{+}$ decays Poblaguev:2002ug , as well as the CHARM experiment CHARM:1985anb . Particularly, rough estimates for potential bounds obtained from charged-pion and kaon decays are also overlaid with dashed lines, for the PIONEER experiment PIONEER:2022yag (blue), kaon factories Goudzovski:2022vbt (green), and the LHC (purple) Altmannshofer:2022ckw . We observe that while these future experiments are estimated to probe mainly the unexcluded parameter region of $c_{ee}/\Lambda\gtrsim 6\times 10^{-4}$ GeV-1, SHiNESS can be sensitive to a unique, currently allowed region $c_{ee}/\Lambda\sim\mathcal{O}(10^{-5})$ GeV-1 for $m_{a}$ between roughly 10 MeV and 120 MeV.
Figure 6: The projected sensitivity of SHiNESS with two years’ operation to the weak-violating electrophilic ALP, displayed in the $(m_{a},c_{ee}/\Lambda)$ plane. The area inside the red curve can be excluded by SHiNESS. The gray area in the background is the currently excluded parameter space, reproduced from Ref. Altmannshofer:2022ckw ; see the text for detail. We also reproduce from Ref. Altmannshofer:2022ckw estimates of potential sensitivities from the PIONEER experiment PIONEER:2022yag (dashed blue), kaon factories Goudzovski:2022vbt (dashed green), as well as a benchmark of the LHC (dashed purple) Altmannshofer:2022ckw . See also the discussion in Sec. 2.2.
@@ 5 Conclusions
In this work, we have investigated the potential of the SHiNESS experiment proposed to be installed at the ESS facility, to search for certain light long-lived particles. Following the SHiNESS proposal Soleti:2023hlr , we focus on the signature of LLPs at SHiNESS as excesses in electron-positron pair events, and selection cuts on the final-state leptons are imposed in order to suppress background events to an acceptable level. Moreover, in order to estimate the cut efficiencies and acceptance of SHiNESS to LLPs, we have performed numerical simulations.
The focus of this work is on LLPs produced in rare decays at rest of $\pi^{+}$ mesons that arise from the 2-GeV proton beam at the facility hitting a tungsten target and lose energy as they propagate. Specifically, we choose two theoretical models predicting light, MeV-scale LLPs, namely, a light bino in the RPV-SUSY, and a weak-violating electrophilic ALP. We consider production modes of the light bino including decays of $\pi^{+}\to e^{+}\tilde{\chi}^{0}_{1}$ and $\mu^{+}\to e^{+}\overset{(-)}{\nu}\tilde{\chi}^{0}_{1}$, where $\mu^{+}$ stems from the $\pi^{+}\to\mu^{+}\nu_{\mu}$ decay at rest and also decays at rest immediately. The light bino then travels and decays inside SHiNESS into an electron-positron pair in association with a neutrino. Three benchmark scenarios, each with two non-vanishing RPV couplings mediating the production and decay of the light bino, respectively, are studied. Included is also one benchmark where a single RPV coupling induces both the production and decay of the light bino. Numerical results are then obtained and presented in the plane of either the production coupling vs. the decay coupling, or the RPV couplings (assumed to be equal) versus the light bino’s mass. Our findings show that SHiNESS is predicted to be able to exclude large, new territories in the parameter space with just two years’ data collection for the considered benchmark scenarios.
For the ALP, we focus on the $\pi^{+}\to e^{+}\,\nu_{e}\,a$ decay for its production and on $a\to e^{+}e^{-}$ for its signal decay. We assume a single, weak-violating term in an effective Lagrangian with the ALP coupled to a pair of electrons, that, after integration by parts, leads to both a helicity-suppressed three-point-vertex term, and a helicity-unsuppressed four-point-vertex term. The latter contribution, involving in the vertex an ALP, an electron, an electron neutrino, as well as a $W$-boson, results in large production rates of the ALP from the $\pi^{+}\to e^{+}\nu_{e}a$ decay. In the numerical results, we find that with solely two years’ operation, SHiNESS can probe a unique, unexcluded part of the parameter space of $c_{ee}/\Lambda\sim\mathcal{O}(10^{-5})$ GeV-1 for $m_{a}$ between 10 MeV and 120 MeV, complementary to other future experiments which would be sensitive to larger values of $c_{ee}/\Lambda$ in a similar range of $m_{a}$.
@@ Acknowledgements
We thank Liangwen Chen, Martin Schürmann, and Stefano Roberto Soleti, for useful discussions. Y.Z. is supported by the National Natural Science Foundation of China under Grant No. 12475106 and the Fundamental Research Funds for the Central Universities under Grant No. JZ2023HGTB0222. W.L. is supported by National Natural Science Foundation of China (Grant No. 12205153).